OpenStudy (anonymous):

Ann and Clark create custom hat bands that people can put around the crowns of their hats/ the finished bands are the lengths shown in the table. The strip of fabric for each band must include the finished length plus an extra 3/4in for the seam. What length of fabric strip is needed to make 4 small bands,5 medium bands, and 3 large bands, including the seam allowance? First, estimate the answer. The use a calculator to find the exact answer. Band size finished lnegth Small: 21 3/8in Medium: 23 1/8in Large: 24 1/4

OpenStudy (pfenn1):

4 x (small band size finished length + 3/4 in) + 5 x (medium band size finished length + 3/4 inch) + 3 x (large band size finished length + 3/4 inch) = total strip length Estimate by rounding. 21 3/8 + 3/4 can be estimated to be around 22 inches 23 1/8 + 3/4 approximately equal to 24, and 24 1/4 + 3/4 = 25 so estimate 4 (22) + 5 (24) + 3 (25)

OpenStudy (anonymous):

so its 25?

OpenStudy (anonymous):

hello? -_-

OpenStudy (anonymous):

88+120+75=

OpenStudy (pfenn1):

Sorry, I had to take my daughter to a class. Yes, 88+120+75=282 for an estimate.

OpenStudy (pfenn1):

Answer 21 3/8 + 3/4 = ? 23 1/8 + 3/4 = ? 24 1/4 + 3/4 = ?